On Wednesday there were no tandem bikes in the shop. There were only bicycles and tricycles. There are a total of 24 seats and 61 wheels in the shop. How many bicycles and how many tricycles are in the shop?
Answer by Edwin McCravy(20059)

On Wednesday there were no tandem bikes in the shop. There were only bicycles and tricycles. There are a total of 24 seats and 61 wheels in the shop. How many bicycles and how many tricycles are in the shop?
Two ways to do it: (1) without using algebra and (2) using algebra.

(1) Without algebra:

24x2=48, 61-48=13 tricycles, 24-13=11 bicycles

Explanation:

There are 24 seats, so there are 24 vehicles.
If they were all bicycles, there would be only
24x2=48 wheels.  However there are 61 wheels, 
so the extra 61-48 or 13 wheels must belong to
the tricycles.  So there are 13 tricycles and 
24-13=11 bicycles.

(2) using algebra:

 B +  T = 24
2B + 3T = 61

Solve the first equation for T

                T = 24 - B

Substitute (24-B) for B in

          2B + 3T = 61
     2B + 3(24-B) = 61
     2B + 72 - 3B = 61
          -B + 72 = 61
               -B = -11
                B = 11   


Substitute B = 11 in

                T = 24 - B
                T = 24 - 11
                T = 13
 
Edwin